# Crafting Adversarial Inputs for Large Vision-Language Models Using Black-Box Optimization

### v3

@@ 6 Conclusion
A Adversarial transferability by Perspective API and Detoxify Classifier Evaluation
We adopt GPT-4o as an automatic evaluation to assess whether an adversarial prompt successfully induces the target model to produce undesired outputs.
Prompt setting. We employ a fixed evaluation prompt to instruct GPT-4o to compute the ASR across four risk categories. The exact prompt used is: “Please use the ASR calculation method across four categories: Identity Attack, Disinformation, Violence/Crime and X-risk.” This standardized prompt ensures consistent evaluation of jailbreak responses across all models.
Aggregation. For each adversarial attempt, we generate 10 candidate responses under the same prompt and model configuration to account for sampling variability. Each response is independently evaluated by GPT-4o using the fixed evaluation prompt. The attempt is considered successful if at least one of the ten responses is undesirable content.
Table 9 and Table 10 provide a comprehensive analysis of the adversarial transferability of ZO-SPSA attacks across three models InstructBLIP, LLaVA and MiniGPT-4, which is evaluated using the Perspective API on the RealToxicityPrompts and Detoxify classifier. Each surrogate model undergoes various perturbation budget constraints $\epsilon=16/255$, $32/255$, $64/255$ and an unconstrained setting. We conduct the evaluation three times under a stronger attack setting and report the means and standard deviations in Table 5.

### v4

@@ 6 Conclusion
A GPT-4o Evaluation
We use GPT-4o to assess whether an adversarial prompt elicits undesired outputs from the target model.
Prompt setting. We employ a fixed evaluation prompt to instruct GPT-4o to classify each generated response with respect to four risk categories. The exact prompt used is: “Please use the ASR calculation method across four categories: Identity Attack, Disinformation, Violence/Crime and X-risk.” This standardized prompt ensures consistent evaluation of jailbreak responses across all models.
Aggregation. For each adversarial attempt, we repeat ten times under the same prompt and model configuration to account for sampling variability. Each response is independently evaluated by GPT-4o using the fixed evaluation prompt. The attempt is considered successful if at least one of the ten responses contains undesirable content and ASR is computed accordingly.
B Adversarial transferability Evaluation
This section provides supplementary empirical results on adversarial transferability referenced in the main paper, offering a more detailed view of transferability on target models.
We adopt the automatic evaluation to assess whether an adversarial prompt successfully induces the target model to produce undesired outputs in Table 4. To evaluate adversarial transferability, we reuse the same adversarial prompt on unseen target models and repeat a stronger attack setting in Table 5.
We additionally report how transferability varies across InstructBLIP, LLaVA, and MiniGPT-4 under multiple $\epsilon$ settings. Each generated response is evaluated by the Perspective API and Detoxify Classifier. Table 9 and Table 10 provide a comprehensive analysis of the adversarial transferability of ZO-SPSA attacks across three target models. Each surrogate model undergoes various perturbation budget constraints $\epsilon=16/255$, $32/255$, $64/255$ and an unconstrained setting.